OpenStudy (anonymous):

The table shows the median home prices in Florida. What is the equation of a trend line that models a relationship between time and home prices? Use the equation to predict the median home price in 2020. Year 1940 1950 1960 1970 1980 1990 2000 Median Price($) $20,200 $39,900 $50,400 $62,800 $89,100 $98,700 $119,000 A. y = 1646.67x + 20,200; $151,900 B. y = 49.4x + 20,200; $119,000 C. y = 49.4x + 20,200; $151,900 D. y = 1646.67x + 20,200; $119,000

OpenStudy (anonymous):

@ajprincess this and 4 more that's it

OpenStudy (ajprincess):

what's ur answr?

OpenStudy (anonymous):

sorry but I dunno this one!

OpenStudy (ajprincess):

hmm np.:)

OpenStudy (anonymous):

yeah, so which one do u think

OpenStudy (ajprincess):

let the equation be y=mx+c nw consider the points (1940, 20200) and (2000, 119000) hope u can find m from this.

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

1646.67

OpenStudy (ajprincess):

yup

OpenStudy (anonymous):

so what's the answer now! A or D

OpenStudy (ajprincess):

median=(2020-1940)*1646.67+20200

OpenStudy (anonymous):

how did u get this formula

OpenStudy (ajprincess):

i too am nt sure. jst a sec. let me figure it out if it is the right one.

OpenStudy (anonymous):

btw the answer is 1747733.6!!!

OpenStudy (anonymous):

so noP

OpenStudy (ajprincess):

check out the answer u got

OpenStudy (ajprincess):

it comes as 151933.6

OpenStudy (anonymous):

oh okay fine

OpenStudy (anonymous):

but how did u get the formula

OpenStudy (anonymous):

oh okay I see what u did

OpenStudy (anonymous):

What are the intercepts of the equation? Graph the equation. -3x + 4y = -12

OpenStudy (ajprincess):

i assumed the equation as y=mx+c.

OpenStudy (anonymous):

C. x-intercept: (4, 0) y-intercept: (0, –3)

OpenStudy (anonymous):

is this right!

OpenStudy (ajprincess):

x-the differnce between the yrs. for the first point (1940-1940=0) so the coordinate of the first pont will be (0,20200), consider the last point (2000-1940,119000)=(60,119000). using these two points i found the equation.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

can u check the other question is this the right graph

OpenStudy (ajprincess):

i am chcking it right nw

OpenStudy (anonymous):

tyt

OpenStudy (ajprincess):

r u sure abt the y-intercept?

OpenStudy (anonymous):

A. x-intercept: (–3, 0) y-intercept: (0, 4) B. x-intercept: (4, 0) y-intercept: (0, –3) C. x-intercept: (4, 0) y-intercept: (0, –3) D. x-intercept: (–3, 0) y-intercept: (0, 4)

OpenStudy (ajprincess):

sorry. made a mistake. ur both intercepts r right:)

OpenStudy (anonymous):

and was the graph right

OpenStudy (ajprincess):

yup right:) the graph for the first question wont be a linear one.

OpenStudy (anonymous):

the question is this What are the intercepts of the equation? Graph the equation. -3x + 4y = -12

OpenStudy (ajprincess):

for that ur answer is right:)

OpenStudy (ajprincess):

I meant the other one. median ques

OpenStudy (anonymous):

this is last question Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12 A. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 B. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (ajprincess):

write the eqn in the form y=mx+c

OpenStudy (ajprincess):

m-slope. c-y-intercept

OpenStudy (anonymous):

9y= 11x-12

OpenStudy (ajprincess):

divide both sides by 9. then u will get the equation in the form y=mx+c

OpenStudy (ajprincess):

have u got it?

OpenStudy (anonymous):

it's C

OpenStudy (anonymous):

slope 11/9 and Y intecept is -4/3

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

okay thnx that's it for now when I will start the rest I will hit u up

OpenStudy (ajprincess):

good work. keep it up:)

OpenStudy (anonymous):

thank u so much u really helped

OpenStudy (ajprincess):

welcome. glad to hear.:)